OpenStudy (anonymous):

need help with math algerbra 2 its tha last question on the test i have been trying to solve it for hours i have been doing this 3 hours can not get this question commonnerfer 0 diana recived 65 points on a project for school she can make changes amd recive 3/10 of the missing points back she can make as many corrections as she wants create a formula for the sum of this geometric series and explain the steps in solving for the maximum grade

OpenStudy (anonymous):

i think its a sigma notation

OpenStudy (anonymous):

no I don't think its sigma notation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

?

OpenStudy (anonymous):

its out of 100 so she can get 35 more

OpenStudy (anonymous):

yes ok

OpenStudy (anonymous):

She lost 35 points She can get 3/10 back 35*3/10=105/10=10.5 points maximum for new grade of 75.5 maximum. converging on 75.5 or 10.5 depending on your viewpoint

OpenStudy (anonymous):

she can get .3 back infintly

OpenStudy (anonymous):

?

OpenStudy (anonymous):

she can make as many corrections as she wants

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so i know it gets closer and closer to 100 but will never reach 100 but i need an equasion

OpenStudy (anonymous):

you can work with him if u want to @Loser66 im not 100% sure im right anyways

OpenStudy (anonymous):

im aproaching the 4 hour mark

OpenStudy (solomonzelman):

I am not sure why you need a geometric series for this problem. Diana received 65 points on a project for school she can make changes and receive 3/10 of the missing points back she can make as many corrections as she wants create a formula for the sum of this geometric series and explain the steps in solving for the maximum grade. She is missing 35 points. She can get 3/10th of them back. 3/10 of 35 is 105/10 which simplifies to, 10, 1/2 or 10.5 So the maximum score can be 65+10.5 which is 75.5.

OpenStudy (anonymous):

it can be done infintly

OpenStudy (solomonzelman):

yeah, I don't think so either loser66, or not really. I can't think strait (or at least can't right now)

OpenStudy (anonymous):

i need a formula thats what i cant get

OpenStudy (anonymous):

how u get the 105/10

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

is this a geometric or arithmetic sequence

OpenStudy (anonymous):

series sorry

OpenStudy (anonymous):

is it geometic or artitmetic

OpenStudy (anonymous):

it needs to be geometic

OpenStudy (anonymous):

did you get it

OpenStudy (anonymous):

ok thank u